the sum of the first and third of three consecutive intergers is 152. fidn the three even integers
Found 2 solutions by checkley77, MathTherapy:
Answer by checkley77(12844)

LET X, X+2 & X+4 BE THE 3 INTEGERS.
X+X+4=152
2X=152-4
2X=148
X=148/2
X=74 THE SMALLEST INTEGER.
74+2=76 THE MIDDLE INTEGER.
74+4=78 THE LARGEST INTEGER.
PROOF:
74+78-152
152=152

Answer by MathTherapy(10858)